OpenStudy (anonymous):

Find the surface area of a rectangular prism that is 16 inches long, 12 inches wide, and 5 inches high.

OpenStudy (calculusfunctions):

Surface Area of a closed rectangular Prism equals to twice the area of the base plus the perimeter of the base multiplied by the height. In your question: Area of the base = (16)(12) = 192 Twice the Area of the base = 2(192) = 384 Perimeter of the base = 2(16) + 2(12) = 56 Surface Area = Twice the Area of the base + the Perimeter of the base x height S.A. = 384 + 56 x 5 ∴ the surface area is 664 square inches.

OpenStudy (anonymous):

woww thank you for explaining it! I totally get it now :) A lot of people just post answers, which isn't very helpful. xD

OpenStudy (calculusfunctions):

Welcome!